OpenStudy (rayman5766):

Solve for x. Express the answer in simplest form. 1/7 = 3/4x - 1

OpenStudy (rayman5766):

@TheRaggedyDoctor help me you left

OpenStudy (trojanpoem):

\[\frac{ 3 }{ 4x -1 } = \frac{ 1 }{ 7 }\] 4x - 1 = 21 4x = 22 x= 5.5 x = 6